OpenStudy (anonymous):

Jerry is experimenting with chemicals in the laboratory. He mixes a solution that is 10% acid with a solution that is 30% acid. How much of the 30% acid solution will be needed to make 40 liters of solution that is 25% acid?

OpenStudy (anonymous):

Let's start with an equation. \[0.10x+0.30(40-x)=0.25*40\] Or \[10x+30*40-30x=25*40\] \[-20x=-5*40\] x=10 liters (amount of 10% solution required) 40-x=30 liters (amount of 30% solution required) You need 30 liters of 30% solution

OpenStudy (anonymous):

thank you mossyfish